OpenStudy (anonymous):

A rectangle has width the same as a side of a square whose perimeter is 20m. the length of the rectangle is 9m. find the perimeter of this rectangle

OpenStudy (e.mccormick):

OK. WHere are you getting stuck on this?

OpenStudy (anonymous):

how to do the actual problem

OpenStudy (anonymous):

wat's ur answer chocies

OpenStudy (e.mccormick):

Well, the first thing you need is the width of the rectangle. So you start with the perimeter of 20m and work backwards from that.

OpenStudy (anonymous):

i dont have answer choices its just a word problem and i have to do it

OpenStudy (e.mccormick):

Now, a square has 4 sides that are the same length. So what does that mean any side is if all 4 add to 20?

OpenStudy (anonymous):

yes the square perimeter is 20m so its 5m per side. is that right

OpenStudy (anonymous):

what i dont understand is what do you mean by start backwards. how do i start the problem

OpenStudy (anonymous):

Yes.

OpenStudy (e.mccormick):

OK. So, you have Rectangle Width = Square Side. Square Side = 5. So what does the rectangle width =?

OpenStudy (e.mccormick):

That is working back from one thing to another.

OpenStudy (anonymous):

so is the square width is 5m

OpenStudy (anonymous):

i mean rectangle width

OpenStudy (e.mccormick):

Yes, the Rectangle Width is 5m. Now they gave you a rectangle length of 9m. So now that you have both the width and the length, should be simple.

OpenStudy (anonymous):

okay so p=2L +2W P=2(9) + 2(5) P=18+10 P=28 IS THIS CORRECT?

OpenStudy (e.mccormick):

That is perfect!

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

can you help me with another question i had

OpenStudy (e.mccormick):

What I emant by backwards is this: They gave you the answer to another parimeter problem. So from that answer, you found some information for this problem. That is the backwards part. Yes, I can help with another.

OpenStudy (anonymous):

yea now i understand i just got confused on how they asked the question and okay one sec

OpenStudy (anonymous):

the second angle of a triangle is twice the first angle.the third angle is 20 degress more than the first angle. find the first angle

OpenStudy (e.mccormick):

Yah, word problems are kind of made that way. The whole deal with them is seeing if you can separate the useful stuff from the extra junk, then apply it. OK. So 3 angles withing a triangle. Do you remember what all the angles of a triangle add up to?

OpenStudy (anonymous):

yes all angles of a triangle add up to 180

OpenStudy (anonymous):

yea sometimes i get it write and sometimes i just get stuck

OpenStudy (e.mccormick):

OK, so lets call first A, second B, and 3rd C so there are less numbers. What they told you is: B=2A C=A+20 And you know A+B+C=180. What they want is A=?

OpenStudy (e.mccormick):

Can you see a way to change A+B+C=180 into A=?

OpenStudy (anonymous):

so if i plug it in A+2A+A+20=180 4A+20=180 4A+20-20=180-20 4A=160 4A/4=160/4 A=40 IS THIS CORRECT

OpenStudy (anonymous):

sorry for caps

OpenStudy (e.mccormick):

Hehe. I did not do the math myself, but that is the process, so I bet you got it. Not hard numbers there.

OpenStudy (e.mccormick):

The key was in gettign rid of first, second, third... which are numbers, and turning them into variables that you could work with. I eman, saying 2nd=2*1st and 3rd=1st+20 is just harder because all the extra numbers.

OpenStudy (anonymous):

yea it was harder thank you for teaching to do it this way it was alot easier with variables

OpenStudy (e.mccormick):

Making them A, B, and C just makes it look more like all the math books. I could have used x, y, z. But, the books like to use A for an angle and a across from it for a side. So if you rad a word probem, work it into the book format that you know and are comfortable with. Then solve it.

OpenStudy (anonymous):

alrighty thank you so much for your help :)

OpenStudy (e.mccormick):

np. Have fun!